# Accreditation Against Limited Adversarial Noise

### v2

@@ III.2.1 Statistical Basis of the New Accreditation Protocol
Before presenting the accreditation protocol, it is first useful to present a purely statistical lemma (Lemma 4) that will later enable the accreditation protocol. This presentation first begins with Def. 9 and Lemma 3.
@@
For any set of real values , =$\mathcal{R}=\big{\{}r_{j}\big{\}}^{|\mathcal{R}|}_{j=1}$, define: $\underline{\mathrm{Avg}\big{(}\mathcal{R}\big{)}}=\dfrac{1}{|\mathcal{R}|}\sum_{j=1}^{|\mathcal{R}|}\bigg{(}r_{j}\bigg{)}.$
@@
Given a set of real values, $\mathcal{R}\subset[(1-\beta)P_{0},(1+\beta)P_{0}]$; if $\beta\in\mathbb{R}$ is known, $P_{0}$ can be bounded as:
$\dfrac{\mathrm{Avg}(\mathcal{R})}{1+\beta}$ $\leq P_{0}\leq\dfrac{\mathrm{Avg}(\mathcal{R})}{1-\beta}.$ (5)
Proof.
Denote =$\mathcal{R}=\{r_{j}|1\leq j\leq|\mathcal{R}|\}$, then:
$\mathrm{Avg}(\mathcal{R})$ $=\dfrac{1}{|\mathcal{R}|}\sum_{j=1}^{|\mathcal{R}|}\bigg{(}r_{j}\bigg{)}.$ (6)
As $\forall r_{j}\in\mathcal{R}$, $r_{j}\in[(1-\beta)P_{0},(1+\beta)P_{0}]$; $\forall r_{j}\in\mathcal{R}$, $(1-\beta)P_{0}\leq r_{j}\leq(1+\beta)P_{0}$. Therefore,
$(1-\beta)P_{0}$ $=\dfrac{1}{|\mathcal{R}|}\sum_{j=1}^{|\mathcal{R}|}\bigg{(}(1-\beta)P_{0}\bigg{)}$ (7)
$\leq\mathrm{Avg}(\mathcal{R})$ (8)
$\leq\dfrac{1}{|\mathcal{R}|}\sum_{j=1}^{|\mathcal{R}|}\bigg{(}(1+\beta)P_{0}\bigg{)}=(1+\beta)P_{0}$ (9)
$\Rightarrow\dfrac{\mathrm{Avg}(\mathcal{R})}{1+\beta}$ $\leq P_{0}\leq\dfrac{\mathrm{Avg}(\mathcal{R})}{1-\beta}.$ (10)
I may then present and prove Lemma 4, which makes use of Lemma 3. Whereas Lemma 3 was concerned with characterizing the interval $\mathcal{R}$, Lemma 4 aims to - using Lemma 3 - place a bound on any value occurring within $\mathcal{R}$ by estimating a bound on $P_{0}$.
Lemma 4.
Given a set of real values, $\mathcal{R}\subset[(1-\beta)P_{0},(1+\beta)P_{0}]$; if $\beta\in[0,1)$ is known, $\forall y\in[(1-\beta)P_{0},(1+\beta)P_{0}]$,
@@
Using Lemma 3,
$P_{0}\leq\dfrac{\mathrm{Avg}(\mathcal{R})}{1-\beta}.$ (12)
$\forall y\in[(1-\beta)P_{0},(1+\beta)P_{0}]$, $y\leq(1+\beta)P_{0}$. Therefore, $\forall y\in[(1-\beta)P_{0},(1+\beta)P_{0}]$,
@@
Using the Taylor series of $\dfrac{1}{1-\beta}$, which holds for all values $\beta$ can take, and neglecting quadratic (or higher) terms:
@@
where I have assumed that the approximation is tight enough to not affect the inequality. ∎
@@ III.2.2 Core Mechanics of the New Accreditation Protocol
Before its formal presentation, in Sec. III.2.3, the core mechanics of the accreditation protocol are presented. This takes the form Lemma 5.
@@
If given a classical polynomial algorithm, $P_{\textit{trap}}^{\prime}$, generating traps corresponding to a fixed circuit, $\mathcal{C}$, and the corresponding $m$, and $k$ (as defined in Sec. III.1); assuming the probability of error in each execution is independent of the outcomes of all preceding executions, that any trap detects any specific error with probability at least $k\in[0,1]$, and all traps are executed via Protocol 1, the probability of error in any execution of any circuit executed within the same singular use of Protocol 1 - such as a target circuit  [^10] As the probability of error in the target would, due to the assumptions made, be within $[(1-\beta)P_{0},(1+\beta)P_{0}]$. - as the traps can be upper bounded by:
@@
using $N_{\mathrm{Tr}}=\bigg{\lceil}\dfrac{2}{\theta^{2}}\ln{\bigg{(}\dfrac{2}{1-\alpha}\bigg{)}}\bigg{\rceil}+1$ traps, where: $\bar{v}$ is the fraction of traps giving an incorrect measurement outcome, $\theta\in\mathbb{R}^{+}$ may be chosen arbitrarily, $\beta$ is as in Protocol 1, and $\alpha$ is the confidence required in the bound.
@@
Let $\mathcal{R}$ be an ordered set where the $j$th element is the probability of error occurring in the $j$th execution of a set of trap executions interspersed among any number of other circuits. In a singular use of Protocol 1. This implies that the probability of error in any specific trap mentioned above - and in any circuit executed within the same single use of Protocol 1 - is within the interval:
$[P_{0}(1-\beta),P_{0}(1+\beta)].$ (16)
Therefore, due to Lemma 4, for any circuit executed in the aforementioned single use of Protocol 1, labelled circuit $j$, the probability of error occurring in its execution, $p_{j}$, is bounded by:
@@
The probability a specific trap returns an incorrect measurement outcome, given error occurs, is, by assumption of having the required traps, lower bounded by $k\in[0,1]$.
Therefore, by also using the independence between circuit executions of whether error occurs, $\mathrm{Avg}(\mathcal{R})$ can be approximated - to within error, $\theta$, with confidence $\alpha$ using Hoeffding’s inequality [50] and $N_{\mathrm{Tr}}=\bigg{\lceil}\dfrac{2}{\theta^{2}}\ln{\bigg{(}\dfrac{2}{1-\alpha}\bigg{)}}\bigg{\rceil}+1$ trap executions.
This approximation of $\mathrm{Avg}(\mathcal{R})$ is simply the fraction of traps returning an incorrect measurement outcome divided by $k$ (to account for the cases where error is not detected). Let $\bar{\nu}$ denote this approximation. Then, with confidence, $\alpha$,
@@ III.2.3 Formal Presentation of the Accreditation Protocol and Proof of its Correctness
The core components of my accreditation protocol have now been constructed and presented - in Sec. III.2.1 and Sec. III.2.2 - so I can now present my accreditation protocol formally, in Algorithm 3. This algorithm is presented as would be used by Alice, presupposing the problem setting in Sec. II.4.
@@
    $\{\mathcal{C}_{j}\}_{j=1}^{N_{l}}$ = $N_{l}$ circuits generated using $P_{\textit{trap}}$ on input $\mathcal{C}$.
@@
    $\mathcal{C}^{\prime}$ = circuit generated using $P_{\textit{targ}}$ on input $\mathcal{C}$.
@@
    Execute every circuit in $\{\mathcal{C}_{j}\}_{j=1}^{N_{l}}\cup\{\mathcal{C}^{\prime}\}$ in a random order using Protocol 1.
@@
The correct functioning of Algorithm 3, as an accreditation protocol, is proven in Theorem 1.
@@
Assuming all sets of circuits executed are executed via Protocol 1, and that the classical algorithms, $P^{\prime}_{\textit{targ}}$ and $P^{\prime}_{\textit{trap}}$, generate trap and target circuits with the features listed in Sec. III.1 (i.e. as in Ref. [23]), then Algorithm 3 provides a sample from a probability distribution and a bound on the ideal-actual variation distance between that distribution and the distribution generated by the measurements in $\mathcal{C}$ when no error occurs (to within accuracy $\theta$, with confidence $\alpha$).
@@
All circuits executed in Algorithm 3 are executed within a single use of Protocol 1. Therefore, if each probability of error in a given trap is independent, then Lemma 5 implies that, as the target circuit is a circuit executed alongside $N_{\mathrm{Tr}}=\bigg{\lceil}\dfrac{2}{\theta^{2}}\ln{\bigg{(}\dfrac{2}{1-\alpha}\bigg{)}}\bigg{\rceil}+1$ traps in a single use of Protocol 1, the probability of error in the target circuit’s execution, $p_{t}$, is upper bounded as:
@@
where $\bar{v}$ is the fraction of traps returning an incorrect measurement result and $k\in[0,1]$ is as defined in Sec. III.1. As mentioned, for Eqn. 19 to hold, it is required that the probability of error in each trap is independent of the outcomes of every prior trap. This is not initially guaranteed as Bob is free to choose probabilities based on previous measurement outcomes. However, due to the changes made to each trap and target in Algorithm 2, to hide the outputs of each trap from Bob with perfect security, Bob does not know the result of any trap or target. Therefore, he cannot make any of his choices based on the results of any prior trap or target  [^11] I additionally note that Bob cannot remove this independence by basing his choice of CPTP maps on error in previous traps. Although Bob knows which CPTP maps are applied to all previous circuits executed, he does not know which errors have occurred due to the stochastic processes involved in the CPTP maps he chose, only the probabilities of each occurring as he doesn’t control any aspect of the CPTP maps’ applications, due to the construction of Protocol 1. Therefore, Bob cannot base the choice of CPTP maps in a trap on the outcome of any stochastic property of previous traps as they are unknown to him.. Hence, each trap can be considered to be independent i.e. the probability of error in each trap or target is the result of Bob’s choices but must remain completely independent of the results of any prior trap or target.
As per Protocol 1, when Bob is given the set of circuits to execute - which in this case would be the traps with the target hidden among them - the single-qubit gates are redacted: therefore, as the targets and the traps only differ in their single-qubit gates (as assumed in Sec. III.1) Bob cannot distinguish traps from targets at any stage in Protocol 3. The redaction additionally allows for the unbreakable encryption of the output of each trap or target, as in Algorithm 2, so Bob cannot locate the target via the outputs either. ∎
Theorem 1 is a proof of the correctness of the protocol developed herein, but the main aim of this paper is to provide Alice with a solution to her predicament outlined in Sec. II.1 and Sec. II.3. To this end, I present another theorem - without proof - that follows directly from Theorem 1 catering to Alice more directly.
Theorem 2.
Protocol 3 allows Alice to get Bob to implement any quantum circuit of her choosing and be able to place a bound on the ideal-actual variation distance (with arbitrary confidence $\alpha$) of the results she receives.
@@ IV Discussion
In this paper I have upgraded the accreditation protocol of Ref. [23] to consider all error to be adversarial. This has necessitated the development of a new model of adversarial error, where Bob is limited according to experimentally extracted rules. These rules - that single-qubit gates are hidden from Bob and that the probability of error in similar circuits executed in quick succession have similar probabilities of error - allow my upgraded accreditation protocol to function almost identically when experiencing adversarial error as the one in Ref. [22] does when the error is IID. Therefore, the desirable qualities the protocol from Ref. [22], such as not requiring lengthening the circuit to accredit, or adding excessively many extra single-qubit gates, or any extra two-qubit gates, or ancilla qubits, or trusting any aspect of a computation, or extra connectivity between qubits are preserved; leading to no diminution in its suitability for near-term usage.
The principle question left open is to what extent can the limits on Bob be relaxed. Relaxing the limitations would have to entail changes to Protocol 1. I believe the redaction of single-qubit gates cannot be eliminated without substantial changes to the trap, as otherwise Bob could distinguish target and trap circuits - although it is not clear trap and target circuits that are indistinguishable to Bob can be constructed without requiring another limitation on Bob.
A slight improvement can be obtained using Ref. [23, Theorem 1], which allows the error in single-qubit gates to be weakly gate dependent: this could be added to my cryptographic model by allowing Bob to define error in the - still redacted - single-qubit gates that depends on the single-qubit gates  [^12] Single-qubit gates would still have to be redacted to prevent Bob identifying the target but he could - perhaps - specify some dependence without seeing the single-qubit gates e.g. “if that gate is an $\hat{X}$ gate…” but the differing CPTP maps - for each single-qubit gate - must differ, in the diamond norm, by at most some small known value. This is a further step towards reality and the restriction on the differences between the gate-dependent error reflects the very small errors in single-qubit gates [32, Fig.5].
Finally I note that, if I take the philosophical position that error that does not affect the state of the qubits, or only adds an overall phase, effectively does not occur and can be considered to have never been applied, I can relax the assumption that Bob does not directly control stochastic processes in the CPTP maps modelling error. This is as the Pauli twirling, accompanied by the random Hadamard gates, make it impossible for Bob to know whether gates he has applied have any affect - which, as mentioned above, can be interpreted as him not being able to decide if he has effectively applied the error. This breaks any ability of Bob to apply error in a way that is dependent on the error in previous circuit executions. Due to this, the error can be considered to be independent between circuit executions.

### v3

@@ III.2.1 Statistical Basis of the New Accreditation Protocol
Before presenting the accreditation protocol, it is first useful to present a purely statistical lemma (Lemma 3) that will later enable the accreditation protocol. This presentation first begins with Def. 14.
@@
For any set of real values, =$\mathcal{R}=\big{\{}r_{j}\in\mathbb{R}^{+}\big{\}}^{|\mathcal{R}|}_{j=1}$, define:
$\underline{\mathrm{Avg}\big{(}\mathcal{R}\big{)}}=\dfrac{1}{|\mathcal{R}|}\sum_{j=1}^{|\mathcal{R}|}\bigg{(}r_{j}\bigg{)}.$ (7)
Later the set $\mathcal{R}$ will denote set of the respective probabilities of error in each trap circuit execution within a single use of Protocol 1 but for the duration of Lemma 3 $\mathcal{R}$ is just considered to contain positive real numbers with no meaning attached to them.
@@
Given a set of positive real values, $\mathcal{R}\subset[(1-\beta)P_{0},(1+\beta)P_{0}]$ (for some $P_{0}\in\mathbb{R}^{+}$); if $\beta\in[0,1)$ is known, $\forall y\in[(1-\beta)P_{0},(1+\beta)P_{0}]$,
@@
Let =$\mathcal{R}=\{r_{j}\in\mathbb{R}^{+}\text{ }|\text{ }1\leq j\leq|\mathcal{R}|\}$. By assumption,
$\forall r_{j}$ $\in\mathcal{R},r_{j}\in[(1-\beta)P_{0},(1+\beta)P_{0}]$ (9)
$\Rightarrow\forall r_{j}$ $\in\mathcal{R},(1-\beta)P_{0}\leq r_{j}.$ (10)
Therefore,
$(1-\beta)P_{0}$ $=\dfrac{1}{|\mathcal{R}|}\sum_{j=1}^{|\mathcal{R}|}\bigg{(}(1-\beta)P_{0}\bigg{)}\leq\dfrac{1}{|\mathcal{R}|}\sum_{j=1}^{|\mathcal{R}|}\bigg{(}r_{j}\bigg{)}=\mathrm{Avg}(\mathcal{R})$ (11)
$\Rightarrow P_{0}$ $\leq\dfrac{\mathrm{Avg}(\mathcal{R})}{1-\beta}.$ (12)
Then, as $\forall y\in[(1-\beta)P_{0},(1+\beta)P_{0}]$, $y\leq(1+\beta)P_{0}$; using Eqn. 12, I conclude that $\forall y\in[(1-\beta)P_{0},(1+\beta)P_{0}]$,
@@
Using the Taylor series of $1/(1-\beta)$ (which converges for all values $\beta$ can take) and neglecting quadratic (or higher) terms:
@@
where I have assumed that the approximation from neglecting higher order terms in the Taylor series is tight enough to not affect the inequality. ∎
@@ III.2.2 Core Mechanics of the New Accreditation Protocol
Before the formal presentation of our new accreditation protocol, in Sec. III.2.3, the core mechanics of the accreditation protocol are presented. This takes the form of Lemma 4.
@@
Given an efficient classical algorithm, $P_{\textit{trap}}^{\prime}$, for generating trap circuits; assuming that:
  1. 1.
    the probability of error in each execution is independent of the outcomes of all preceding executions
  2. 2.
    any trap detects any specific error, by outputting specific measurement outcomes, with probability at least $k\in[0,1]$
  3. 3.
    all trap circuits are executed via a single use of Protocol 1
the probability of error of any execution of a circuit where all error is Pauli twirl, during the same single use of Protocol 1 as the trap circuit executions can be upper bounded by:
@@
using $N_{\mathrm{Tr}}=\bigg{\lceil}\dfrac{2}{\theta^{2}}\ln{\bigg{(}\dfrac{2}{1-\alpha}\bigg{)}}\bigg{\rceil}+1$ trap circuit executions, where:
$\bullet$ $\bar{v}\in\mathbb{Q}$ is the fraction of trap circuit executions giving an incorrect measurement outcome,
$\bullet$ $\theta\in\mathbb{R}^{+}$ may be chosen arbitrarily,
$\bullet$ $\beta\in\mathbb{R}^{+}$ is as in Protocol 1,
$\bullet$ $\alpha\in[0,1]$ is the confidence required of the bound in Eqn. 15.
@@
Let $\mathcal{R}$ be an ordered set where the $j$th element is the probability of error occurring in the $j$th execution of a set of trap executions, interspersed among any number of other circuits, in a singular use of Protocol 1.
This implies that the probability of error in any specific trap (as generated by $P_{\textit{trap}}^{\prime}$) – and in any circuit executed within the same use of Protocol 1 as those trap circuit executions (such as target circuits generated by $P_{\textit{targ}}^{\prime}$) – is within an interval that may be written as:
$[P_{0}(1-\beta),P_{0}(1+\beta)],$ (16)
for some $\beta,P_{0}\in\mathbb{R}^{+}$. Therefore, due to Lemma 3, for any circuit, labeled circuit $j$, executed in the aforementioned single use of Protocol 1, along the trap circuit executions, the probability of error occurring in its execution, $p_{j}$, is bounded by:
@@
The probability a specific trap returns an flags/detects that an error occurs by returning an incorrect measurement outcome, given error occurs, is (as assumed in the lemma statement) lower bounded by $k\in[0,1]$.
Therefore, by also using the independence between circuit executions of whether error occurs, $\mathrm{Avg}(\mathcal{R})$ can be approximated, by checking if the trap circuit executions detect error, – to within additive error, $\theta$, with confidence $\alpha$ – using Hoeffding’s inequality [52]. This only requires that $|\mathcal{R}|\geq N_{\mathrm{Tr}}=\bigg{\lceil}\dfrac{2}{\theta^{2}}\ln{\bigg{(}\dfrac{2}{1-\alpha}\bigg{)}}\bigg{\rceil}+1$, to provide enough trap circuit executions that the approximation of $\mathrm{Avg}(\mathcal{R})$ is within the required error, $\theta$.
Let $\bar{\nu}$ denote the experimentally obtained approximation of the probability a randomly selected (from $\mathcal{R}$) trap returns an output implying error has occurred (i.e. the fraction of trap circuit executions returning an “incorrect” measurement outcome in a single execution of Protocol 1).
This approximation of $\mathrm{Avg}(\mathcal{R})$ is simply $\bar{\nu}$ divided by $k$ (to account for the cases where error occurs but is not detected). Then, with confidence, $\alpha$, the probability of error occurring in the execution of the circuit with index $j$ is bounded as:
@@ III.2.3 Formal Presentation of the Accreditation Protocol and Proof of its Correctness
The core components of my accreditation protocol have now been constructed and presented – in Sec. III.2.1 and Sec. III.2.2 – so I can now present my accreditation protocol formally, in Algorithm 3. This algorithm is presented as would be used by Alice, presupposing the problem setting in Sec. II.4.
@@
    Choose a random integer $N_{tt}$ less than $10\cdot N_{l}$.
@@
    $\{\mathcal{C}_{j}\}_{j=1}^{N_{l}+N_{tt}}$ = $N_{l}+N_{tt}$ circuits generated using $P^{\prime}_{\textit{trap}}$ on input $\mathcal{C}$.
@@
    $\mathcal{C}^{\prime}$ = circuit generated using $P^{\prime}_{\textit{targ}}$ on input $\mathcal{C}$.
@@
    Execute every circuit in $\{\mathcal{C}_{j}\}_{j=1}^{N_{l}}\cup\{\mathcal{C}^{\prime}\}$ in a random order (i.e. disregarding the indexing used herein) using Protocol 1.
  6. 6.
@@
The correct functioning of Algorithm 3, as an accreditation protocol, in the adapted/modified problems setting is proven in Theorem 1.
@@
In the adapted cryptographic setting (as defined in Sec. II); given efficient classical algorithms, $P^{\prime}_{\textit{targ}}$ and $P^{\prime}_{\textit{trap}}$, that generate trap and target circuits, respectively (as in Sec. III.1); Protocol 3 allows Alice to execute a circuit equivalent (in output distribution, when no error occurs) to any sampBQP circuit, $\mathcal{C}$, and obtain a bound on the ideal-actual variation distance of that execution (to within arbitrary accuracy $\theta\in\mathbb{R}^{+}$, with arbitrary confidence, $\alpha\in[0,1]$).
@@
All circuits executed in Protocol 3 are executed within a single use of Protocol 1 (in step 5 of Protocol 3). Therefore, if each probability of error in a given trap is independent, then Lemma 4 implies that, as the target circuit is a circuit executed alongside $N_{\mathrm{Tr}}=\bigg{\lceil}\dfrac{2}{\theta^{2}}\ln{\bigg{(}\dfrac{2}{1-\alpha}\bigg{)}}\bigg{\rceil}+1$ trap circuit executions in a single use of Protocol 1, the probability of error in the target circuit’s execution, $p_{t}$, is upper bounded as:
@@
where $\bar{v}$ is the fraction of trap circuit executions returning an incorrect measurement result and $k\in[0,1]$ is as defined in Sec. III.1.
As mentioned, for Eqn. 19 to hold, it is required that the probability of error in each trap is independent of the outcomes of every prior trap. This is not initially guaranteed as Bob is free to choose probabilities based on previous measurement outcomes.
However, due to the changes made to each trap and target in Algorithm 2, to hide the outputs of each trap from Bob with perfect security, Bob does not know the result of any trap or target and, as the limitation on Bob that means the probability of error due to the error he applies cannot be deterministic, Bob – as he is not told about the outcomes of the random processes in the error he applies – is unaware of what errors have been applied in each circuit execution.
Additionally, due to the $N_{tt}\in\mathbb{N}_{0}$ extra trap circuits executed during Protocol 1, Bob does not know what he has done to the previous circuits as he has no idea which are “decoy” circuit executions and will be discarded (so he cannot know what error he has applied to the “true” circuit executions). In fact, as Bob is not told $\theta,\alpha$, or $N_{tt}$, he cannot even calculate the various probabilities that he has applied each possible combination of errors – or lack thereof – to to the “true” trap circuit executions.
Therefore, Bob cannot make any of his choices based on the results of any prior trap or target execution or the error that occurs in them. Hence, each trap execution can be considered to be independent i.e. the probability of error in each trap or target is the result of Bob’s choices but must remain completely independent of the results of any prior trap or target.
As per Protocol 1, when Bob is shown the set of circuits to execute (which in this case would be the trap circuit with the target circuit hidden among them), by Robert, the single-qubit gates are redacted. Therefore, as the target circuits and the trap circuits only differ in their single-qubit gates (as assumed in Sec. III.1) Bob cannot distinguish trap circuit executions from target circuit executions at any stage in Protocol 3. The redaction additionally allows for the unbreakable encryption of the output of each trap or target, as in Algorithm 2, so Bob cannot locate the target via the outputs either.
With the probability of error in any target circuit, $\mathcal{C}_{\text{targ}}$, executed within the same single use of Protocol 1 bounded – as in Eqn. 19 – a bound on the ideal-actual variation distance of the execution of that target circuit, $\nu\big{[}\tilde{\mathcal{C}}_{\text{targ}}\big{]}$, can be obtained. By following Ref. [23, Appendix Sec. 1], I derive that the ideal-actual variation distance of the execution of that target circuit is upper bounded by the probability of error occurring in that execution. Therefore,
$\nu\big{[}\tilde{\mathcal{C}}_{\text{targ}}\big{]}\leq\dfrac{1+2\beta}{k}\big{(}\bar{v}+\theta\big{)}.$ (20)
@@ IV Discussion
In this paper I have upgraded the accreditation protocol of Ref. [23] to consider all noise/error to be adversarial. This has necessitated the development of a new model of adversarial error (i.e. the adapted cryptographic setting presented in Sec. II), where Bob is limited according to experimentally derived rules. These rules (that single-qubit gates are hidden from Bob and that the probability of error in similar circuits executed in quick succession have similar probabilities of error), encapsulated in Protocol 1, allow my upgraded accreditation protocol to function almost identically when experiencing adversarial error as the one in Ref. [22] does when the error is IID. Therefore, the desirable qualities the protocol from Ref. [22], such as not requiring lengthening the circuit to accredit, or adding excessively many extra single-qubit gates, or any extra two-qubit gates, or ancilla qubits, or trusting any aspect of a computation, or extra connectivity between qubits are preserved; leading to no diminution in its suitability for near-term usage.
The principle question left open by this paper is to what extent the limits on Bob can be relaxed. Relaxing the limitations would have to entail changes to Protocol 1. I believe the redaction of single-qubit gates cannot be eliminated without substantial changes to the trap, as otherwise Bob could distinguish target and trap circuits, although it is not clear trap and target circuits that are indistinguishable to Bob can be constructed without requiring another limitation on Bob.
A slight improvement can be obtained using Ref. [23, Theorem 1], which allows the error in single-qubit gates to be weakly gate dependent: this could be added to my cryptographic model by allowing Bob to define error in the – still redacted – single-qubit gates that depends on the single-qubit gates [^11] Single-qubit gates would still have to be redacted to prevent Bob identifying the target but he could, perhaps, specify some dependence without seeing the single-qubit gates e.g. “if that gate is an $\hat{X}$ gate…”. but the differing CPTP maps – for each single-qubit gate – must differ, in the diamond norm, by at most some small known value. This is a further step towards reality and the restriction on the differences between the gate-dependent error reflects the very small errors in single-qubit gates [34, Fig.5].